OpenStudy (anonymous):

Which equation represents the line passing through the points (3, 2) and (−9, 6)? x − 3y = 9 x + 3y = 9 3x − y = −9 3x + y = 9

OpenStudy (anonymous):

Plug in the variables (3, 2) and (−9, 6). x − 3y = 9 (3)-3y = 9 -3y = 6 y = -2 Not the equation. x + 3y = 9 (3) + 3y = 9 3y = 6 y = 2 (-9) + 3y = 9 3y = 18 y = 6 This is the equation. 3x − y = −9 3(3) - y = -9 y = 18 Not the equation. 3x + y = 9 3(3) + y = 9 y = 0 Not the equation.